OpenStudy (anonymous):

how do i evaluate this expression? 27^ 1/3 so far i did 3√ 27 but i don't know the next step

OpenStudy (anonymous):

27=3^3 so (3^3)^1/3= 3^1=3

OpenStudy (anonymous):

or\[\sqrt[3]{3^3}=3\]

OpenStudy (anonymous):

the first approach uses the property\[\left( b^u \right)^v=b^{uv}\]

OpenStudy (anonymous):

very helpful (: thank you so much

OpenStudy (anonymous):

ur welcome :})

OpenStudy (anonymous):

one more question how do you make that square root symbol?